# Toward Deep Drum Source Separation

### v1

@@ 1 Introduction
A revised and updated version of the present manuscript has been submitted to Pattern Recognition Letters on Sept. 26, 2023.
@@
Clips Duration (h) ${}^{\star}$ Drum kits Real/Synth Human Mixture Transcription Isolated stems
@@
ENST-Drums (Gillet and Richard, 2006) $456$ $3.75$ ${}^{\star}$$20/8^{\star}$ $3$ R $\checkmark$ Stereo TXT $\checkmark$
@@ 5.2 Results
First, none of the methods exhibits a noticeable drop in performance when evaluated on the four held-out drum kits (marked with ${}^{\dagger}$ in Table 3), suggesting that all three methods can generalize in the face of unseen timbral characteristics.
@@
Table 3: nSDR computed on StemGMD Eval Session. The six drum kits denoted by the superscript ${}^{\star}$ were included in the training dataset, whereas the four drum kits denoted by ${}^{\dagger}$ were held out for evaluation.
@@
Brooklyn East Bay Heavy Portland Retro Rock SoCal ${}^{\dagger}$ Detroit ${}^{\dagger}$ Motown ${}^{\dagger}$ ${}^{\dagger}$ All

### v2

@@ Abstract
Peer-Reviewed Version
A peer-reviewed version of the manuscript has been published in Pattern Recognition Letters and is available at 10.1016/j.patrec.2024.04.026. Please, cite this work as: A. I. Mezza, R. Giampiccolo, A. Bernardini, and A. Sarti, "Toward Deep Drum Source Separation," Pattern Recognition Letters, 2024, doi: 10.1016/j.patrec.2024.04.026.
@@ 1 Introduction
Clips Duration (h) Classes/Mics Drum kits Real/Synth Human Mixture Transcription Isolated stems
@@
ENST-Drums (Gillet and Richard, 2006) $456$ $3.75$ $20/8$ $3$ R $\checkmark$ Stereo TXT $\checkmark$
@@ 5.2 Results
First, none of the methods exhibits a noticeable drop in performance when evaluated on the four held-out drum kits (marked with in Table 3), suggesting that all three methods can generalize in the face of unseen timbral characteristics.
@@
Table 3: nSDR computed on StemGMD Eval Session. The six drum kits denoted by the superscript  were included in the training dataset, whereas the four drum kits denoted by  were held out for evaluation.
@@
Brooklyn East Bay Heavy Portland Retro Rock SoCal Bluebird Detroit Garage Motown Rev. Roots All